J&L Steel Company

All Notices [87] rss feed
Categories
<<  <  1 2 3 4 5  ...  >  >> 
Notices  Posted Category
Bulletin #85 9/4/2015 9/8/2015 Site Update
On September 3, 2015 ER continued to gather and remove soil along the bank of the Little River, smoothing out the slope of the riverbank. Soil clean or contaimianted is being placed in the appropriate staging area. Dust controls were in place with the use of the watertruck to suppress particualt dust onsite.

Full-time site activities will shutdown the week of September 7, 2015 and will resume on September 14, 2015. ER deployed additional sorbent boom along the lower portion of the Little River to mitgate oil seeps coming off the bank.

Future activities at the Site include conducting additional LIF drilling the week of September 21 and will plan on conducting sonic core drilling the week of September 28 and October 5, 2015.


Bulletin #84 9/1/2015 9/1/2015 Site Update

ER is working on removing the soil and rock along the bank of the Little River. Most of the material is clean and is being placed in the staging area to the southwest. ER is also working on replacing sorbent boom along the Little River. RST 3 is compiling the test pit information from last week for EPA.



Bulletin #83 8/27/2015 8/27/2015 Site Update

EPA ERT / SERAS continued with the subsurface investigation at the J&L Steel Site using the LIF drilling method to measure PAHs in the soil. Today the LIF rig advanced westward past the Red Building. Drilling proceeded slower today as refusal was encountered. The highest fluorescence reading so far was detected about 100 feet northeast of the Red Building.

EPA ERT conducted additional investigative activities on the west side of the site. Six test pits were excavated to investigate groundwater near the high density polyethylene (HDPE) barrier wall. Oil was not observed on the west end of the barrier wall. Oil was observed in several test pits excavated against the HDPE barrier wall on both sides with heavy sheening and strong odors. It was evident based on these observations that oil contaminated groundwater is making its way under the HDPE barrier and migrating to the Little River via the seeps observed along the bank which also bubble up from the bottom of the river. It does not seem that it is making its way around the barrier wall. This hasn’t been observed in the whole liner yet. This will be something to look at in the future.

ER continued to load more soil away from the bank of the Little River.



Bulletin #82 8/26/2015 8/26/2015 Site Update

EPA ERT/ SERAS are directing the investigation of subsurface oil at the J&L Steel Site. Laser Induced Fluorescence (LIF) technology is being used to assess poly-aromatic hydrocarbons (PAHs) in the soil at various points across the Site. Zebra Tech is using a Geo Probe GP44 model to advance a probe tip which emits an UV-laser, which measures the PAH signature directly in front of the window. Drilling began on August 24, 2015 with points being drilled towards the northeast side of the Site moving west. An electric conductivity (EC) meter was combined with the LIF probe to measure soil conductivity.

Normandeau Environmental continued with vibracore drilling downstream of the culvert bridge. Several cores where hand drilled close to the bank directly downstream. Soils being brought up to the surface were analyzed at the field station which were initially screened using a TVA-1000. Soils were identified and then were analyzed for TPH analysis onsite.

ER continued to remove soil piles along the bank of the Little River which were placed during drilling activities last week. Soil that was contaminated was placed in the contaminated staging area to the northeast. Clean soil was placed in the staging area to the southwest.



Bulletin #81 8/19/2015 8/19/2015 Site Update

EPA continued with the soil investigation of the Little River bottom. Normandeau Environmental continued to conduct vibracore drilling operations. They resumed their track from yesterday finishing the borings downstream along the center of the channel near the culvert bridge. The drillers then focused on collecting borings towards the north bank, placing borings every 100 ft. then 200 ft further upstream. The drillers completed the borings along the north bank and began placing borings along the southern bank. The drillers encountered frequent refusal while drilling along the southern bank. A total of 16 cores were extracted today.

The mobile lab which is being utilized by EPA ERT / SERAS was moved near the red building / ERRS storage building. Cores extracted from the river were field screened for VOCs using a TVA-1000, using the PID and FID. Soils were identified and then placed in jars where they were analyzed in the mobile lab for TPH analysis. Some samples will be shipped for analytical confirmatory analysis.

ERRS continued to remove soil along the bank of the Little River segregating the clean soil from the contaminated and placing in the appropriate staging area. Sections of the sorbent boom along the bank of the Little River had to be repositioned.Dust controls were enforced throughout the day. The water truck sprayed down the main access roads.

Town officials from the Clifton-Fine office and the NY State DEC were onsite to take a look at progress at the J&L Steel Site. A tour of the site was given showing the areas of concern at the Site as well as the drilling activities.



Bulletin #80 8/18/2015 8/18/2015 Site Update

EPA ERT, SERAS, and Weston RST 3 conducted an investigation of the Little River by analyzing the soil along the bottom of the Little River. Normandeau Environmental Consultants are the drilling contractors onsite. The vibracore drill method was used to collect cores samples from the river. The core is advanced using a vibrating push mechanism which advances a plastic sleeve into the river bottom extracting the core sample. Fourteen cores were extracted today along the middle of the Little River every 100 feet from upstream down to about 200 feet from the culvert bridge. The cores were then extracted and field analyzed using a TVA-1000 using the PID and FID. Samples were collected which will be analyzed using the mobile lab onsite.

 ER continued to remove clean soil along the bank of the Little River which was placed in the staging area. Dust controls were enforced due to high temperatures and dry conditions generating dust.



Bulletin #79 8/17/2015 8/17/2015 Site Update

ER is removing soil piles along the bank of the Little River and moving to the clean fill to the appropriate staging area. Dust control measures were in place due to dry conditions which could generate particulate dust. The water truck was used to wet the roadways.

EPA ERT, Lockheed Martin/ SERAS and Weston RST3 are going to be conducting the investigation of the sediment in the Little River. Normandeau Environmental Consultants arrived at the Site and will be conducting the collection of core samples using the vibracore drilling method. Their boat was placed in the water today and will begin drilling tomorrow.



Bulletin #78 8/12/2015 8/12/2015 Site Update

ER continued to work by the bank of the Little River. Rock which has been stockpiled temporarily close to the bank was being moved to the clean staging area. Large rocks will be turned into rip rap. ER also continued cleaning up the bank of the Little River removing rocks along the bank and some vegetation. RST 3 monitored water quality near the downstream culvert bridge where oil sheen has been observed flowing more frequently as soil is being disturbed. Additional sorbent boom was replaced starting near the culvert bridge and moving upstream. Additional boom will be placed in this area to contain any oil migrating downstream. There are about 4 seeps that have been observed which is originating from subsurface migration of oil towards the stream which has moved past the pre-existing polyethylene barrier wall.



Bulletin #77 8/11/2015 8/11/2015 Site Update

ER utilized an excavator to remove rocks and vegetation along the bank of the Little River. RST 3 observed oil sheen on the surface of the Little River seeping from the bank about 100 yards upstream from culvert bridge. ER began replacing sorbent boom along the riverbank with additional boom being placed in areas where oil sheen is the heaviest. A bulldozer was also used to smooth out the bank surface.  Verizon arrives to connect the data/phone lines and remove the old lines from the old trailer. Now the old trailer is ready to be removed.  



Bulletin #76 8/6/2015 8/6/2015 Site Update

ER continues to excavate along the bank of the Little River. Oil contaminated water was observed seeping out as the soil which was mixed with large rocks was being removed. Oil sorbent boom was replaced downstream to contain any oil the migrated past the culvert bridge. Dust controls were in place today with the water truck watering down the road surfaces.



Bulletin #75 8/5/2015 8/5/2015 Site Update

ER continues to excavate the soil from a natural ledge which extends from the bank of the Little River near the downstream culvert bridge. Most of the soil and rocks being excavated are contaminated with oil which is being placed in the appropriate staging area to the northeast of the Site. Plastic liner was placed underneath the soil and then covered at the end of the day to protect from precipitation and runoff. Due to dry weather and little soil moisture, dust control measures were activated with the use of the water truck. However, the pump used to fill the truck stopped working. A new pump will arrive tomorrow to assist in dust control measures. Vehicle speeds were kept at 10 mph throughout the day to minmize dust.



Bulletin #74 8/4/2015 8/4/2015 Site Update

ER conducted activities along the bank of the Little River today near the downstream culvert bridge. This included the removal of the natural ledge which diverts the river flow downstream. ER worked on opening up this ledge so water can pass through while leaving some of vegetative barrier in place to contain any oil seepage. The bank surface was also smoothed out as well. RST 3 monitored any oil release as a result of work being performed. Some oil sheen was observed in areas where sediment was disturbed.



Bulletin #73 7/30/2015 7/30/2015 Site Update

On 7/29/2015, ER began removing the railroad tracks near the former scale house, north of the new office trailer. There was standing water observed underneath the scale platform but no oil was detected. Once the tracks were removed soil was placed to fill in the area and smoothed to grade.



Bulletin #72 7/29/2015 7/29/2015 Site Update

ER continues to clear the area west of the new office trailer location. During clearing activities the operator noticed an area that had some hydrocarbon odors directly south of the new trailer. RST3 field screened the area with a MultiRae with no detections. Dust controls were in place yesterday as temperatures reached above 90o F. Workers made sure to stay hydrated and took frequent breaks. ER began cleaning out a 500-gallon poly tank which will be used to pump out contaminated groundwater which was generated from the groundwater investigation activities.



Bulletin #71 7/28/2015 7/28/2015 Site Update

The front gate of the Site was found damaged yesterday morning. Repairs were made during most of the morning yesterday. ER continued to perform site maintenance activities at the Site including grading the site and smoothing access roads. Temperatures reached 90o F yesterday with very little soil moisture on the ground. Dust control measures were in place which include watering down the access roads. Heat stress and hydration were the safety concerns for yesterday and the rest of the week. The area directly southwest of the new office trailer is being regraded. This was the former scrap metal debris pile which has been moved closer to the clean fill staging area.